# DoubleAgents: Exploring Mechanisms of Building Trust with Proactive AI

### v1

@@ 4.2. Interactive Chat Interface
Upon clicking “Start Planning,” Lee is directed to the main Interactive Chat Interface (Figure 2(B)), where he can view the full history of system interactions and monitor the planning and execution progress carried out by the core ReAct agents. All plans, actions, and emails generated by the system are accessible through this interface. Each system update appears as an interactive pop-up window (Figure 2(C) shows a pop-up plan item), enabling Lee to inspect and confirm the system output in a controllable and traceable way. The interface also supports natural language queries, giving Lee full control over what to view. For example, he can type “What did Prof. B reply in the last email?” to retrieve the relevant message history directly within the chat. This chat functionality is powered by GPT-4o, with prompt details provided in the Appendix A.4.
4.3. Coordination Agent
4.3.1. Progress Understanding
When Lee initiates the planning process, the coordination agent assesses the overall state of the problem by generating a concise summary of the progress so far. To do that, it first reviews the current state of the scheduling process by analyzing the problem context, including speakers’ info, email logs from previous days, and the history of plans and actions previously approved by Lee. Based on this information, the coordination agent generates a summary of the current status, indicating which days have been scheduled, what actions have been taken, and the most recent contact date for each speaker to help monitor responses and follow-ups. The prompt used in generating the summary is illustrated in the Appendix A.1.1.
4.3.2. Policy Selection
To design a set of general policies applicable to a broad range of planning problems, five experts were recruited to brainstorm and define a framework of rules tailored to address common scheduling challenges. These challenges include: (1) individuals being reluctant to share their full availability, (2) ambiguous or unclear responses, (3) the need for multiple follow-ups to elicit responses, (4) resolving scheduling conflicts between two speakers, and (5) initiating the process by broadly reaching out to potential participants. Based on these identified challenges, we developed six initial ”seed policies” as a core set of guidelines for the system to reference and select from when generating plans. Each policy is designed to address a specific challenge, with two policies dedicated to handling non-responsive participants: one focused on strategies to cope with the lack of response (follow-up frequency) and another providing a framework for when to give up on follow-ups.
In Lee’s example, since it is the first day of scheduling, the progress summary informs the system that no speaker availabilities have been collected yet. As a result, the coordination agent invokes the policy selection tool, which identifies the appropriate policies based on the current problem state. Drawing from the list of ”seed policies”, the system selects: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” The selected policy is then highlighted in the Policy Panel (as shown in Figure 2(A)), providing Lee with the rationale to generate the plan. The policy acts as the planning guideline, informing the agent’s next steps. Details of the prompts used for the policy selection module are provided in the Appendix A.1.2.
@@ 4.3.4. Action Generation
To operationalize the plan, the coordination agent then generates a set of concrete actions to be executed (see Figure 4(b)). Specifically, it proposes sending individual emails to each speaker, with the stated goal: “Request for availability for Seminar Slots (Sep 5, Oct 3, Nov 7, Dec 5).” These proposed actions are presented to Lee for review. He examines each action and its corresponding goal, deems them appropriate, and clicks Approve to proceed. Importantly, Lee retains full control over the suggested actions: he may edit the goals, remove actions that are misaligned, or instruct the system to regenerate the set based on revised input. In general, there are three types of actions: (1) drafting emails to selected respondents, (2) sending emails to specified recipients using their provided addresses, and (3) awaiting responses from certain respondents. The detailed prompts used for plan and action generation are provided in the Appendix A.1.3.
@@ 4.3.6. Implicit User Preference Learning
DoubleAgents supports implicit learning of user preferences by observing user-initiated edits to system-generated plans, actions, or emails. For instance, if the coordination agent proposes to reach out to Student X to ask for more availability, Lee may choose to remove the action “Send initial email to Student X”—perhaps because Student X’s prior responses have been vague or unhelpful. The system treats such edits as preference signals. To act on these signals, the coordination agent summarizes the edit into a concise statement, logs it in a user edit history, and incorporates it into the problem context. This updated context is then referenced in subsequent planning steps, enabling the agent to evolve its behavior in alignment with the user’s preferences over time. To streamline the summarization process, we leverage GPT-4o. Details of the prompt used for generating summaries of user edits across various content types can be found in Appendix A.1.5. The user-initiated edits, the corresponding system-generated drafts, and the generated summaries of the edits are all preserved in the user edit history. This comprehensive history serves as an evolving knowledge base, directly influencing the system’s generation of future plans, actions, and emails, and ensuring that it remains responsive to the nuances of user behavior.
@@ 4.4. Simulated Respondent Agent
To construct a realistic simulation environment for Lee, we draw inspiration from (Park et al., 2023) and employ LLMs to simulate each respondent agent, generating actions that align with their assigned personas. Concretely, upon receiving Lee’s emails, each simulated respondent agent first reasons over its persona and then generates an action that adheres strictly to those characteristics. For instance, the agent representing Prof. A opts not to respond on the same day (shown in Figure 1), as his persona specifies: “You are very busy and reply very late. Normally, it takes you three follow-up emails to respond.” In contrast, Prof. B’s persona indicates she is prompt in communication, leading her agent to reply immediately. Given that Prof.B’s persona also specifies preferences, such as “prefers presenting over Zoom”, the generated action will explicitly request a Zoom presentation option. The prompt for speaker agent, both for generating actions and drafting email responses, is detailed in the Appendix A.3. To execute the selected actions, the simulated respondent agents compose and send email responses, which are then received by Lee on his end. Figure 5(a) shows an example response received by Lee as shown in the interface. These responses advance the overall scheduling process and enable Lee to continue coordinating subsequent steps. Similar to the coordination agent, we adopt GPT-o3 as the backbone of each respondent agent to perform reasoning and generate plans and actions. GPT-4o is used when invoking the sendEmail() tool to generate email responses.
@@ 4.5. Edge Case Detector
Within a day, Lee receives an email from Prof. B stating: “Thanks for reaching out, Lee! I was wondering if there might be an option to present via Zoom instead of in person? That would provide more flexibility for me.” As this type of request had not been previously addressed by any existing policy, representing an edge case, the edge case detector issues a flag to prompt clarification from Lee. Through the Interactive Chat Interface, Lee sees the following system message: “Prof. B requests a Zoom option due to circumstances preventing in-person attendance. Should I confirm Zoom availability for their seminar or insist on in-person participation?” Preferring in-person interactions for seminars, Lee responds with the instruction: “Tell Prof. B we prefer in-person due to the potentially large audience.” The coordination agent then processes this clarification and generates a new action to inform Prof. B accordingly. Lee reviews the suggested action, finds it appropriate, and clicks Approve. An email draft is subsequently generated; after confirming the message aligns with his intent, Lee clicks ’Approve & Send’ to dispatch the response to Prof. B. In general, the edge case detector acts as a tool for the coordination agent to check if the message sent from the respondent agent can be addressed with the existing set of policies in order to make the next plans. When the detector identifies an edge case, the coordination agent escalates the issue to the user by presenting a clarification question in the interactive chat interface, prompting the user to provide guidance on how to proceed. The edge case detector is implemented with GPT-4o (see prompts in the Appendix A.2).
@@ 4.6. Dashboard Visualization
To support user understanding and foster trust, DoubleAgents provides clear visual feedback on the task state and system progress throughout the scheduling process. We implement a Flask-based dashboard interface designed to make the planning workflow interactive, legible, and actionable. The dashboard is composed of several layered visual components, shown in Figure 2(D), (E), and (F), which we describe in detail below:
@@
At the end of each simulated day, the system generates a daily summary that highlights key events and decisions. Specifically, we use GPT-4o to generate the daily summaries, with prompt details provided in the Appendix A.5. As shown in Figure 2(B), one example summary is presented within the Interactive Chat Interface and reads: “Invitations for the HCI seminar were sent to Profs. A, B, and Students X, Y. Prof. B and Student Y expressed interest, with scheduling pending. Both Prof. B and Student X inquired about Zoom options instead of in-person attendance.” After reviewing the summary, Lee confirms the day’s progress and clicks Continue to Next Day to proceed. This iterative, day-by-day process continues until all seminar slots are successfully assigned.
@@ 5.1.1. Experiment settings
We evaluated the effectiveness of our approach by comparing it against three baseline variants, each using a reduced context: (1) speaker-only context, which includes only information about speaker responsiveness; (2) slot-only context, which includes only slot-level availability (e.g., filled vs. unfilled); and (3) combined speaker and slot context, which merges the two but omits higher-level summaries. Prompt templates for all conditions are provided in Appendix B. To evaluate the performance of each approach, we constructed a test suite of 30 examples, all sourced from real user study sessions and curated to reflect a diverse set of scheduling scenarios. For each example, we treat the set of policies approved by users during the study as the ground truth for evaluation. All evaluations were conducted using the o4-mini model with a default temperature of 1. We report the average F1 score across three independent runs for each experimental setting.
@@ 5.1.2. Results and Findings
Prompting
Techniques
F1 Score
@@
5.2. Edge Case Detection Experiments
5.2.1. Experiment settings
Edge case detection involves identifying whether a respondent’s message presents a scenario that falls outside the coverage of existing coordination policies. For example, a speaker requesting to present via Zoom, when no policy has addressed virtual presentations, should be flagged to the user for clarification and further instruction. To enable reliable detection of such edge cases, we adopt a few-shot prompting strategy with access to existing policies (denoted as ”few-shot w/ policy”). We compare this approach against two baselines: (1) zero-shot prompting without any demonstration examples, and (2) few-shot prompting without access to policy references (denoted as ”few-shot w/o policy”). Concrete prompt templates for each setup are provided in Appendix B.2. To evaluate performance, we constructed a test suite of 100 examples, manually validated by the lead authors to ensure coverage and quality. The construction methodology for this test suite is detailed in Appendix B.2. All experiments were conducted using GPT-4o, with temperature set to 0.7 and a maximum token limit of 100. We report the average accuracy across three independent runs for each condition.
5.2.2. Results and Findings
Prompting Techniques
Accuracy
@@ 5.2.2. Results and Findings
As shown in Table 2, our few-shot prompting with policy method significantly outperforms both baselines, with statistically significant improvements across evaluation runs. These results demonstrate the effectiveness of incorporating policy context into the detection process, enabling the model to first analyze existing policies before determining whether a given case falls outside their scope. Based on these findings, DoubleAgents adopts the policy-aware few-shot approach as its default mechanism for edge case detection. This ensures that novel or ambiguous cases are reliably flagged to users, enabling timely clarification and maintaining a high-quality user experience.
@@ 6. User Study
6.1. Goal and Setup
To investigate the user-perceived effectiveness of DoubleAgents’ features and the dynamics of trust in human-AI collaboration, we recruited 10 participants to use DoubleAgents across two separate sessions. Our goal was to assess users’ perceived system usefulness, their reliance on the systems, and their trust and comfort in using AI to automate tasks had shifted. During each session, participants worked in a simulated environment to schedule a seminar with four dates and four speakers, while the system gradually adapted to their preferences. Participants completed a questionnaire assessing their trust in 15 different system features at three points: before the first session, after the first session, and after the second session. This allowed us to examine whether and when their trust and perceptions of the system changed over time.

### v2

@@ 3.2. Interactive Chat Interface
Upon clicking “Start Planning,” Lee is directed to the main Interactive Chat Interface (Figure 3(B)), where he can view the full history of system interactions and monitor the planning and execution progress carried out by the core ReAct agents. All plans, actions, and emails generated by the system are accessible through this interface. Each system update appears as an interactive pop-up window (Figure 3(C) shows a pop-up plan item), enabling Lee to inspect and confirm the system output in a controllable and traceable way. The interface also supports natural language queries, giving Lee full control over what to view. For example, he can type “What did Prof. B reply in the last email?” to retrieve the relevant message history directly within the chat. This chat functionality is powered by GPT-4o, with prompt details provided in the Appendix C.4.
@@
3.3. Coordination Agent
3.3.1. Progress Understanding
When Lee initiates the planning process, the coordination agent assesses the overall state of the problem by generating a concise summary of the progress so far. To do that, it first reviews the current state of the scheduling process by analyzing the problem context, including speakers’ info, email logs from previous days, and the history of plans and actions previously approved by Lee. Based on this information, the coordination agent generates a summary of the current status, indicating which days have been scheduled, what actions have been taken, and the most recent contact date for each speaker to help monitor responses and follow-ups. The prompt used in generating the summary is illustrated in the Appendix C.1.1.
3.3.2. Policy Selection
To design a set of general policies applicable to a broad range of planning problems, the lead authors brainstorm and define rules tailored to address common scheduling challenges. These challenges include: (1) individuals being reluctant to share their full availability, (2) ambiguous or unclear responses, (3) the need for multiple follow-ups to elicit responses, (4) resolving scheduling conflicts between two speakers, and (5) initiating the process by broadly reaching out to potential participants. Based on these identified challenges, we developed six initial ”seed policies” as a core set of guidelines for the system to reference and select from when generating plans. Each policy is designed to address a specific challenge, with two policies dedicated to handling non-responsive participants: one focused on strategies to cope with the lack of response (follow-up frequency) and another providing a framework for when to give up on follow-ups.
In Lee’s example, since it is the first day of scheduling, the progress summary informs the system that no speaker availabilities have been collected yet. As a result, the coordination agent invokes the policy selection tool, which identifies the appropriate policies based on the current problem state. Drawing from the list of ”seed policies”, the system selects: “If we are just beginning and have not yet contacted any speakers for availability, then ask each speaker for their availability across all slots.” The selected policy is then highlighted in the Policy Panel (as shown in Figure 3(A)), providing Lee with the rationale to generate the plan. The policy acts as the planning guideline, informing the agent’s next steps. Details of the prompts used for the policy selection module are provided in the Appendix C.1.2.
@@ 3.3.4. Action Generation
To operationalize the plan, the coordination agent then generates a set of concrete actions to be executed (see Figure 5(b)). Specifically, it proposes sending individual emails to each speaker, with the stated goal: “Request for availability for Seminar Slots (Sep 5, Oct 3, Nov 7, Dec 5).” These proposed actions are presented to Lee for review. He examines each action and its corresponding goal, deems them appropriate, and clicks Approve to proceed. Importantly, Lee retains full control over the suggested actions: he may edit the goals, remove actions that are misaligned, or instruct the system to regenerate the set based on revised input. In general, there are three types of actions: (1) drafting emails to selected respondents, (2) sending emails to specified recipients using their provided addresses, and (3) awaiting responses from certain respondents. The detailed prompts used for plan and action generation are provided in the Appendix C.1.3.
@@ 3.3.6. Implicit User Preference Learning
DoubleAgents supports implicit learning of user preferences by observing user-initiated edits to system-generated plans, actions, or emails. For instance, if the coordination agent proposes to reach out to Student X to ask for more availability, Lee may choose to remove the action “Send initial email to Student X”—perhaps because Student X’s prior responses have been vague or unhelpful. The system treats such edits as preference signals. To act on these signals, the coordination agent summarizes the edit into a concise statement, logs it in a user edit history, and incorporates it into the problem context. This updated context is then referenced in subsequent planning steps, enabling the agent to evolve its behavior in alignment with the user’s preferences over time. To streamline the summarization process, we leverage GPT-4o. Details of the prompt used for generating summaries of user edits across various content types can be found in Appendix C.1.5. The user-initiated edits, the corresponding system-generated drafts, and the generated summaries of the edits are all preserved in the user edit history. This comprehensive history serves as an evolving knowledge base, directly influencing the system’s generation of future plans, actions, and emails, and ensuring that it remains responsive to the nuances of user behavior.
@@ 3.4. Simulated Respondent Agent
To construct a realistic simulation environment for Lee, we draw inspiration from (Park et al., 2023) and employ LLMs to simulate each respondent agent, generating actions that align with their assigned personas. Concretely, upon receiving Lee’s emails, each simulated respondent agent first reasons over its persona and then generates an action that adheres strictly to those characteristics. For instance, the agent representing Prof. A opts not to respond on the same day (shown in Figure 2), as his persona specifies: “You are very busy and reply very late. Normally, it takes you three follow-up emails to respond.” In contrast, Prof. B’s persona indicates she is prompt in communication, leading her agent to reply immediately. Given that Prof.B’s persona also specifies preferences, such as “prefers presenting over Zoom”, the generated action will explicitly request a Zoom presentation option. The prompt for speaker agent, both for generating actions and drafting email responses, is detailed in the Appendix C.3. To execute the selected actions, the simulated respondent agents compose and send email responses, which are then received by Lee on his end. Figure 6(a) shows an example response received by Lee as shown in the interface. These responses advance the overall scheduling process and enable Lee to continue coordinating subsequent steps. Similar to the coordination agent, we adopt GPT-o3 as the backbone of each respondent agent to perform reasoning and generate plans and actions. GPT-4o is used when invoking the sendEmail() tool to generate email responses.
@@ 3.5. Edge Case Detector
Within a day, Lee receives an email from Prof. B stating: “Thanks for reaching out, Lee! I was wondering if there might be an option to present via Zoom instead of in person? That would provide more flexibility for me.” As this type of request had not been previously addressed by any existing policy, representing an edge case, the edge case detector issues a flag to prompt clarification from Lee. Through the Interactive Chat Interface, Lee sees the following system message: “Prof. B requests a Zoom option due to circumstances preventing in-person attendance. Should I confirm Zoom availability for their seminar or insist on in-person participation?” Preferring in-person interactions for seminars, Lee responds with the instruction: “Tell Prof. B we prefer in-person due to the potentially large audience.” The coordination agent then processes this clarification and generates a new action to inform Prof. B accordingly. Lee reviews the suggested action, finds it appropriate, and clicks Approve. An email draft is subsequently generated; after confirming the message aligns with his intent, Lee clicks ’Approve & Send’ to dispatch the response to Prof. B. Lee’s clarification is recorded as a stop hook and used to refine the policy set, explicitly specifying that Zoom is not preferred when speakers request remote participation. In this way, user interventions during edge cases are converted into reusable system policies. In general, the edge case detector acts as a tool for the coordination agent to check if the message sent from the respondent agent can be addressed with the existing set of policies in order to make the next plans. When the detector identifies an edge case, the coordination agent escalates the issue to the user by presenting a clarification question in the interactive chat interface, prompting the user to provide guidance on how to proceed. The edge case detector is implemented with GPT-4o (see prompts in the Appendix C.2).
@@ 3.6. Dashboard Visualization
To support user understanding and foster alignment, DoubleAgents provides clear visual feedback on the task state and system progress throughout the scheduling process. We implement a Flask-based dashboard interface designed to make the planning workflow interactive, legible, and actionable. The dashboard is composed of several layered visual components, shown in Figure 3(D), (E), and (F), which we describe in detail below:
@@
At the end of each simulated day, the system generates a daily summary that highlights key events and decisions. Specifically, we use GPT-4o to generate the daily summaries, with prompt details provided in the Appendix C.5. As shown in Figure 3(B), one example summary is presented within the Interactive Chat Interface and reads: “Invitations for the HCI seminar were sent to Profs. A, B, and Students X, Y. Prof. B and Student Y expressed interest, with scheduling pending. Both Prof. B and Student X inquired about Zoom options instead of in-person attendance.” After reviewing the summary, Lee confirms the day’s progress and clicks Continue to Next Day to proceed. This iterative, day-by-day process continues until all seminar slots are successfully assigned.
@@ 4.1.1. Experiment Settings
We evaluated the effectiveness of our approach by comparing it against three baseline variants, each using a reduced context: (1) speaker-only context, which includes only information about speaker responsiveness; (2) slot-only context, which includes only slot-level availability (e.g., filled vs. unfilled); and (3) combined speaker and slot context, which merges the two but omits higher-level summaries. Prompt templates for all conditions are provided in Appendix D. To evaluate the performance of each approach, we constructed a test suite of 30 examples, all sourced from real user study sessions and curated to reflect a diverse set of scheduling scenarios. For each example, we treat the set of policies approved by users during the study as the ground truth for evaluation. All evaluations were conducted using the o4-mini model with a default temperature of 1. We report the average F1 score across three independent runs for each experimental setting.
@@ 4.1.2. Results and Findings
Prompting Techniques F1 Score
@@
Prompting Techniques Accuracy
@@
4.2. Edge Case Detection Experiments
4.2.1. Experiment Settings
Edge case detection involves identifying whether a respondent’s message presents a scenario that falls outside the coverage of existing coordination policies. For example, a speaker requesting to present via Zoom, when no policy has addressed virtual presentations, should be flagged to the user for clarification and further instruction. To enable reliable detection of such edge cases, we adopt a few-shot prompting strategy with access to existing policies (denoted as ”few-shot w/ policy”). We compare this approach against two baselines: (1) zero-shot prompting without any demonstration examples, and (2) few-shot prompting without access to policy references (denoted as ”few-shot w/o policy”). Concrete prompt templates for each setup are provided in Appendix D.2. To evaluate performance, we constructed a test suite of 100 examples, manually validated by the lead authors to ensure coverage and quality. The construction methodology for this test suite is detailed in Appendix D.2. All experiments were conducted using GPT-4o, with temperature set to 0.7 and a maximum token limit of 100. We report the average accuracy across three independent runs for each condition.
4.2.2. Results and Findings
As shown in Table 2, our few-shot prompting with policy method significantly outperforms both baselines, with statistically significant improvements across evaluation runs. These results demonstrate the effectiveness of incorporating policy context into the detection process, enabling the model to first analyze existing policies before determining whether a given case falls outside their scope. Together, DoubleAgents adopts the policy-aware few-shot approach as its default mechanism for edge case detection. This ensures that novel or ambiguous cases are reliably flagged to users, enabling timely clarification and maintaining a high-quality user experience.
@@ 5. User Study
5.1. Goal and Overview
To investigate how users engaged with DoubleAgents’ features and how interactive simulation shaped their use of agentic behavior, we recruited 10 participants to use DoubleAgents across two separate sessions. Our goal was to assess users’ perceived system usefulness, their understanding and reliance on the system, and how their understanding and calibration of agent behavior evolved over time. During each session, participants worked in a simulated environment to schedule a seminar with four dates and four speakers, while the system gradually adapted to participants’ values and preferences via their policies, email templates, and stop hooks. Participants completed two questionnaires. First, an AI comfortableness survey measured trust and alignment needs for proactive actions and comfort with automating planning and execution, administered before Session 1, after Session 1, and after Session 2. Second, a System perception survey captured UX and perceptions of 15 system features after Session 1 and after Session 2, enabling us to track how repeated use shifted alignment strategies, delegation boundaries, and reliance. Detailed study procedures and data analysis methods are provided in Appendix A.